# Random Scaling of Emergent Capabilities

### v4

@@ 3.1 Data and Setup
We extend our investigation to LMs by testing multiple choice question answering, an emergent natural language understanding task (Srivastava et al., 2023; Snell et al., 2024). We test whether the performance clusters observed in emergent synthetic tasks also manifest in LMs. Specifically, we hypothesize that emergent scaling curves in LMs express underlying bimodal (or multimodal) performance distributions.
@@
Our procedure can be regarded as continued pretraining: we mimic the model’s original pretraining objective with a dataset of English language sequences. We hypothesize that at the emergence point for this task, performance distributions will also be bimodal across random seeds. During pretraining, a large diverse corpus encourages models to acquire various capabilities. While sufficiently large models may learn all such capabilities, smaller models have limited capacity, requiring these capabilities to compete (Merrill et al., 2023). This competition, influenced by initialization and data order, leads to varying outcomes across random seeds, forming performance clusters for specific benchmarks.
@@
Task: We test LMs on MMLU Hendrycks et al. (2021), a multiple-choice question-answering benchmark. Strong performance on MMLU requires two key abilities: (1) natural language reasoning with domain-specific knowledge and (2) producing correct answers in the required format. It is this latter ability which leads to emergent trends (Srivastava et al., 2023; Hu and Frank, 2024).
@@
Model: We use the Qwen2.5 family of base models Yang et al. (2024). To introduce randomness during continued pretraining, we reinitialize the final attention layer and the subsequent LM head. We perform full-parameter continued pretraining on these reinitialized models.
@@
Data: To ensure that the multiple choice formatting circuits compete with other tasks during continued pretraining, we create a diverse dataset by mixing the English news subset of C4 Raffel et al. (2023) with the official MMLU auxiliary training data. The C4 news data reinforces general language modeling ability, while the MMLU data focuses on multiple-choice question answering. We vary the proportion of MMLU in this mixture to control the target task’s data size in addition to model size, as both are common control variables in the scaling laws literature.
@@
Training: We continue pretraining the Qwen2.5-0.5B and Qwen2.5-1.5B models on our C4-MMLU mixes, training 80 reinitializations on each data mixture ratio. We train Qwen2.5-0.5B models for 2 epochs and Qwen2.5-1.5B models for 5 epochs to ensure convergence on the MMLU validation set. We use a learning rate of 1e-5 with linear decay scheduler.
We focus our analysis on experiment results from Qwen2.5-0.5B. In Appendix E, we include results from Qwen2.5-1.5B (Appendix 17) as well as a second task CommonSenseQA (Talmor et al., 2019) (Appendix E.2).
3.2 Emergence Across Data Compositions
We first examine how data composition influences the bimodal capabilities of Qwen2.5-0.5B. Specifically, as the training data contains more examples of a target breakthrough task, it changes the task’s performance distribution in ways that mirror the effect of scale.
In Figure 7 (left), when trained on a data mix containing only 5% MMLU examples and 95% random sequences from C4 Raffel et al. (2023), most models (out of 80 seeds) achieve near 0% performance on the MMLU test set. According to the existing literature, this failure stems from a failure to process the multiple-choice format Hu and Frank (2024). As the proportion of MMLU samples increases, bimodality emerges when MMLU composes 10% of the training data, where models form two distinct performance clusters. In one cluster, continually pretrained models still fail to follow the correct format, while in the other, models achieve performance at or above the random guess baseline of 25%. This second cluster contains models that consistently respond with valid multiple choice options; those that outperform the random baseline have even learned to compose the multiple choice format with their world knowledge when selecting an answer. Finally, with sufficient data ($20\%$ MMLU), models across all 80 seeds consistently outperform the 25% random baseline.
3.3 Emergence Across Model Scales
@@
Like in the synthetic setting, LM model size affects the clustered performance distribution. Using a training data mixture with 10% MMLU auxiliary training samples, we continually pretrain Qwen2.5-1.5B on the same 80 seeds and report MMLU test accuracy in Figure 7 (right). While models built on the smaller Qwen2.5-0.5B form two distinct performance clusters, those built on the larger Qwen2.5-1.5B consistently outperform the random baseline. In fact, larger models reliably acquire MMLU capability when trained on the same dataset that produces highly bimodal variation at smaller scales. This conclusion recalls the literature on scaling laws, which argues that smaller models require more training examples to match the performance of larger models (Rosenfeld et al., 2019; Kaplan et al., 2020). Appendix E.2 shows similar results on the Commonsense QA (Talmor et al., 2019) task.
@@
Figure 8: NLL loss distributions on the MMLU test set. Each distribution is provided as a histogram and as a smoothed KDE. Even with a continuous metric (NLL), we still observe multi-modal distribution across random seeds.
@@
Figure 9: KDE approximation of LM training loss on 10% MMLU mixture. Approximation of the distribution across 80 random seeds is unimodal, smooth, and nearly symmetrical with median close to mean. By contrast, Figure 8 depicts MMLU loss for LMs trained on the same mix as irregularly clustered with multiple modes.
@@ 3.4 Is MMLU emergence a mirage?
Is multiple choice performance only bimodal because accuracy is thresholded? As in our synthetic setting (Section 2.4), we must address the effect of the discontinuous metric. In Figure 8, we find that continuous negative log likelihood (NLL) loss metrics remain highly multimodal. The 10% MMLU ratio leads to the most even split between successes and failure in accuracy, and accordingly to clear bimodality in NLL loss, mirroring the link we observed on synthetic tasks. Notably, training data with a 7.5% MMLU ratio produces a performance distribution with a small “success” cluster peaked around NLL=3.4. The 10% MMLU ratio produces a performance distribution with a larger success cluster, also peaking around NLL=3.4. Therefore, the continuous metrics remain roughly bimodal—as is especially clear in the 10% mix setting, which is most evenly split between success and failure.
@@
While recent arguments against emergence (Schaeffer et al., 2024) claim that discontinuities disappear with continuous loss metrics, Figure 8 reveals the opposite effect in random variation—NLL loss reveals clusters that are concealed by accuracy metrics, specifically among low-performance models like those trained on the 5% MMLU mix. (To easily compare these metrics, see Figure 13.)
@@
Is this random clustering particular to the emergent task? As an alternative hypothesis, we might expect all tasks—including training loss—to have similarly multimodal NLL distributions, when approximated with our KDE settings. To the contrary, the pretraining loss produces a smooth, unimodal, and almost symmetric KDE distribution (Figure 9). We therefore conclude that multimodality is linked to emergent tasks—not merely a downstream effect of in-distribution training noise.
@@ 4 Discussion
Our work explores the evolution of random variation in model performance across scales, bringing a more nuanced perspective on emergent capabilities. The mode of our performance distributions sharply improves at a certain model scale, similar to emergent capabilities in the literature, but we attribute these sudden jumps to gradual improvements in the random distribution. Furthermore, we show that bimodality can emerge before the mode—or most seeds—exhibit a breakthrough (Section 2.5); the transition from the initial unimodal distribution to a bimodal one is sudden.
@@
Random variation: Model performance can be sensitive to stochastic aspects of the training process like random initialization and training data order. Prior studies have documented performance differences across various stress test sets (D’Amour et al., 2022; Naik et al., 2018). Most relevant to us, the high variability of length generalization is well-documented (Zhou et al., 2024b; a). More generally, out-of-distribution behavior like compositional rules (McCoy et al., 2019) or associative biases (Sellam et al., 2022) often exhibit extreme variation compared to in-distribution loss. Such differences persist throughout training, not just at the final checkpoint (Zhou et al., 2020). Dodge et al. (2020) compared the impacts of weight initialization and data ordering, concluding that both contribute equally to variation in performance. Existing work has also found model runs can cluster in shortcut learning (Juneja et al., 2023) and in training dynamics (Qin et al., 2024; Hu et al., 2023), hinting at multimodal variation. Our work connects these random clustering effects to the phenomenon of emergence at scale.
Length generalization and compositionality with transformers: Transformer models cannot always generalize to longer sequences than they are trained on (Anil et al., 2022; Deletang et al., 2023; Gontier et al., 2020; Hupkes et al., 2020; Schwarzschild et al., 2021; Zhang et al., 2022). Various proposals have, consequently, aimed to improve length generalization. Some methods focus on alternative positional encodings (Shaw et al., 2018; Press et al., 2022; Su et al., 2024; Kazemnejad et al., 2024; Jelassi et al., 2024). Others modify the dataset format by adding scratchpad or Chain-of-Thought (Anil et al., 2022) or by incorporating padding and index hints (Jelassi et al., 2023; Zhou et al., 2024a). Regarding random variation, Zhou et al. (2024a) and Zhou et al. (2024b) provide evidence of variability in length generalization across random seeds, which we further investigate and analyze across a range of model scales.
@@
Emergent abilities of LMs: In LLMs, emergent abilities are behaviors that arise unexpectedly as models are scaled up in size or trained on larger datasets (Hestness et al., 2017; Rosenfeld et al., 2019; Brown et al., 2020; Kaplan et al., 2020). These abilities are characterized by unpredictable and abrupt performance improvements on specific benchmarks at certain scales (Wei et al., 2022; Ganguli et al., 2022; Srivastava et al., 2023). Understanding the conditions and mechanisms underlying emergence is a key area of research. Although recent studies suggest that some breakthroughs may stem from the choice of evaluation metrics rather than fundamental changes in model behavior with increased scale (Srivastava et al., 2023; Schaeffer et al., 2024), some breakthrough capabilities remain emergent. Snell et al. (2024) found that some scales exhibit earlier emergence if finetuned explicitly on an emergent task, suggesting that smaller models may have the capacity for that task but are limited by its scarcity in the training corpus. Similarly, we show that emergent capabilities can arise from multimodal random variation using synthetic length generalization tasks as a case study.
@@
Depth versus Width Scaling: Although scaling laws offer a smooth extrapolation for model performance, downstream performance can vary depending on architecture shape and not just model size (Tay et al., 2022). For compositional tasks, deeper models often generalize better up to a certain point, but for a fixed compute budget, it may be more advantageous to train a shallower, wider model (Petty et al., 2024). Various works have proposed explanations for the role of width versus depth in scaling behavior; for instance, Edelman et al. (2024) show that increasing network width offers more ‘parallel queries’ over randomized subnetworks which learn sparse features more efficiently. Levine et al. (2020) use a border rank argument to establish a width-dependent depth threshold, beyond which additional depth yields diminishing returns. In this work, we specifically investigate how independently scaling width and depth influences the random variation distribution in compositional tasks. In one surprising result, we document a regime in which increasing width damages model performance but increasing depth improves model performance. In a research community increasingly concerned with emergent capabilities, our findings should inspire further study of how emergent tasks respond to architectural hyperparameter tradeoffs.
@@ Acknowledgments
Reproducibility Statement
Upon publication, we will release the code for both synthetic and LLM experiments, including code, and data. Detailed descriptions of model, data and training are provided in Appendix A, and Section 3.
@@ Appendix A Experimental details on synthetic tasks
Below we provide more details when training decoder-only language models on the count and addition—with index hints—tasks. Hyperparameters are largely adapted from [Zhou et al., 2024a]. We train all of our models to convergence on the train distribution. Our experiments were run on an internal cluster and all model scales can be run on a single 40GB A100 GPU with gradient accumulation; for count, runs can finish within 2 hours and for addition, runs can finish within 6 hours.
Count task: For all of our training runs, we fix the vocabulary size to 150. For evaluation, we compute the exact match (EM) accuracy across all consecutive subsequences of the test length.
  • Model scales: As mentioned in Section 2.1, we scale up our models by fixing width and scaling depth and fixing depth and scaling width. The precise parameters for each variation are as follows. For our fixed depth experiments, we fix the network depth to 4 layers and vary width by taking hidden dimensions $\{64,128,256,384,512,640,768,1024\}$. The head dimension is fixed to 64. For our fixed width experiments, we fix the hidden dimension to be 512 and vary the depth from $\{1,2,4,6,8\}$ layers.
@@
Reverse Order Addition with Index Hints: For evaluation, we compute the exact match (EM) accuracy across 500 batches of 128 examples each.

### v5

@@ 3.1 Data and Setup
To avoid the expense of repeatedly training large LMs from scratch, we simulate independent runs by reinitializing the upper layer of pretrained LMs before continuing to train them. Our continued training dataset, a mix of C4 news data and MMLU training data, is insufficient to recover the base model’s full MMLU performance. We hypothesize that for compute scales near the emergence threshold, performance distributions will be bimodal across random seeds.
@@
During pretraining, a large diverse corpus encourages models to acquire various capabilities. While sufficiently large models may learn all capabilities, smaller models have limited capacity, requiring capabilities to compete (Merrill et al., 2023). This competition, influenced by initialization and data order, leads to varying outcomes across random seeds, forming performance clusters.
@@
Task: We test LMs on MMLU (Hendrycks et al., 2021), a multiple-choice question-answering benchmark. Strong MMLU performance requires (1) natural language reasoning with domain knowledge and (2) producing answers in the required format. The latter drives emergent trends (Srivastava et al., 2023; Hu and Frank, 2024).
@@
Model: We use the Qwen2.5 family of base models (Yang et al., 2024). To introduce randomness, we reinitialize the final attention layer and LM head, then perform full-parameter continued pretraining.
@@
Data: We mix the C4 news subset (Raffel et al., 2023) with MMLU training data to ensure the multiple choice formatting circuits compete with general language modeling. We vary the MMLU proportion to control task-specific data size.
@@
Training: We continue pretraining Qwen2.5-0.5B and Qwen2.5-1.5B on C4-MMLU mixes, training 80 reinitializations per data mixture ratio. We train for 2 epochs (0.5B) and 5 epochs (1.5B) with learning rate 1e-5 and linear decay.
(a) Qwen2.5-0.5B mix results.
(b) Qwen2.5: 10% MMLU data.
Figure 5: Performance on MMLU. All models have their last layer reinitialized randomly and then are trained on a mix of C4 news and MMLU, with varying proportions of MMLU. (a) With insufficient MMLU training data, performance is trivial; as data increases, performance becomes bimodal across seeds. (b) For a fixed data mix, smaller models show bimodal performance while larger models consistently relearn the task, albeit below original base levels.
3.2 Experiment Results
Emergence Across Data Compositions.
When training data contains more MMLU examples, the Qwen2.5-0.5B performance distribution improves in ways that mirror the effect of scale. In Figure 5 (left), when trained on 5% MMLU and 95% C4 (Raffel et al., 2023), most models (out of 80 seeds) achieve near 0% MMLU accuracy, failing to process the multiple-choice format. At 10% MMLU training data, bimodality emerges: one cluster remains near 0% accuracy (format failure), while the other reaches $\geq$25% accuracy (random baseline). This second cluster contains models that consistently produce valid multiple choice responses; those exceeding 25% have learned to compose format following with some recovered world knowledge. Due to the limited scale and quality of our continued pretraining data, fully recovering the base model’s MMLU capability is challenging. At $20\%$ MMLU, all 80 seeds consistently perform around or above the random baseline. Qwen2.5-1.5B shows the same bimodality (Figure 19), but the larger model recovers similar capabilities at much lower data mixture ratios.
Emergence Across Model Scales.
@@ 3.2 Experiment Results
Model size affects the clustered MMLU performance distribution similarly to the synthetic settings in Sec. 2. Using 10% MMLU training data, we continually pretrain Qwen after reinitializing with 80 different seeds (Figure 5, right). While Qwen2.5-0.5B forms two distinct performance clusters, Qwen2.5-1.5B consistently performs around or above the random baseline. Larger models reliably acquire MMLU capability on the same dataset that produces highly bimodal variation at smaller scales, consistent with scaling laws: smaller models require more training examples to match larger model performance (Rosenfeld et al., 2019; Kaplan et al., 2020).
Figure 6: Question formation performance across data compositions. Left: Mean and median OOD accuracy across 80 seeds shows emergence at 10% hierarchical-rule inducing data. Right-4: Histograms reveal bimodal distributions at 10%, 90% and 100% hierarchical data. Models either cluster at either 0% (linear generalization) or 100% (hierarchical generalization).
Figure 7: Question formation performance across model scales. Training on 90% hierarchical-rule inducing data, 3-layer models show unimodal distribution around 70% OOD accuracy, while 6-layer models exhibit bimodality with most seeds achieving 100% (hierarchical generalization) and some reaching 0% (linear generalization). Each dot represents one seed.
@@
Figure 8: NLL loss distributions. Even with a continuous metric (NLL), we still observe multi-modal distributions in performance across random seeds. NLL is measured on: Left: MMLU for Qwen2.5-0.5B trained on 10% MMLU mixture. Right: OOD QF for LM pretrained on 90% hierarchical mixture.
@@
Figure 9: KDE of LM training loss. The distribution across 80 seeds is unimodal and smooth. By contrast, Fig. 9 depicts MMLU loss as irregularly clustered.
@@ 3.3 Is MMLU emergence a mirage?
Is multiple choice performance only bimodal because accuracy is thresholded? As in our synthetic setting (Sec. 2.4), we examine continuous metrics. In Fig. 9, negative log likelihood (NLL) loss remains bimodal. For the 10% MMLU training data case, we confirm this bimodality is statistically significant ($p<0.001$) for significant using Hartigan’s Dip Test (Hartigan and Hartigan, 1985), which measures the maximum difference between the empirical distribution and the best-fitting unimodal distribution. The bimodality observed in NLL also mirrors our synthetic algorithmic task findings.
@@
While recent work (Srivastava et al., 2023; Schaeffer et al., 2024) argues that discontinuities disappear with continuous metrics, Fig. 19 and Fig. 19 reveals the exact opposite: NLL loss exposes clusters concealed by accuracy metrics, particularly among low-performance models trained on 5% MMLU (Fig. 19) and among larger models (Fig. 19).
@@
Is this clustering specific to emergent tasks? Fig. 9 suggests so. Pretraining loss produces a smooth, unimodal, nearly symmetric distribution, contrasting sharply with MMLU’s multimodality. We therefore conclude that multimodality is linked to emergent tasks. It is not merely a downstream effect of in-distribution training noise.
4 Grammatical Generalization in LMs
We next examine English syntax acquisition during LM training. Question formation (QF) is a canonical benchmark (Mueller and Linzen, 2023; McCoy et al., 2020b, a) for grammatical rule learning. In this setting, the ambiguous training data supports two possible competing rules: (i) a hierarchical syntactic rule (correct for English) or (ii) a superficial linear heuristic (like an n-gram language model). On the disambiguating OOD test data, the hierarchical rule yields  100% accuracy, the linear rule yields  0%, and any score in between indicates inconsistent rule application.
Prior work has documented performance clustering in QF across random seeds (Qin et al., 2024; Ahuja et al., 2024; Murty et al., 2023), with evidence that data composition and model scale influence which rule is learned (Qin et al., 2024; Ahuja et al., 2024). This controlled setting allows us to study bimodal distributions of competing viable solutions when training a simple language model from scratch.
4.1 Data and Setup
[…]
Model: Decoder-only Transformers trained from scratch on causal language modeling. Following prior work (Qin et al., 2024; Murty et al., 2023), we vary depth (3-6 layers) with 8 heads per layer and 512-dimensional embeddings.
Data: Synthetic English sentences generated via Context-Free Grammars (McCoy et al., 2018). Based on Qin et al. (2024), we vary the proportion of hierarchical-inducing versus linear-inducing sentences in the training data, paralleling our MMLU data mixture experiments.
Training: We train 80 random seeds per configuration. We use Adam optimizer with learning rate $1\times 10^{-4}$ and linear decay and train models for 600K steps. We vary (1) model size and (2) proportion of hierarchy-inducing training examples, to study how these scale factors influence the bimodal performance distribution of hierarchical generalization.
4.2 Experiment Results
Emergence Across Data Compositions.
Fig. 6 (left) shows mean OOD accuracy exhibits sharp emergence at 10% hierarchical data, jumping from near 0% to 80%. However, histograms (right 4 panels) reveal that the population mean masks the underlying bimodality: at 10%, seeds split between 0% (linear rule) and 80-100% (hierarchical rule). Varying data composition shifts which rule models learn. More hierarchical-inducing sentences increase the probability of learning the hierarchical rule, but do not eliminate the competing solution even at 100% hierarchical-inducing data. Apparently, even our most hierarchical-inducing training set offers insufficient support to provide consistently “successful” runs, possibly because it lacks diversity.
Emergence Across Model Scales.
Fixing the data mixture at 90% hierarchical sentences, we now examine how model capacity affects the performance distribution. At 3 layers (Fig. 7), performance is unimodal around 70%, with few seeds reaching 100%—these smaller models lack capacity to learn either systematic syntactic rule. At 6 layers, clear bimodality emerges: most seeds achieve 100% (hierarchical generalization) while some reach 0% (linear generalization). Although our largest scale models tend to consistently apply systematic rules, we find different seeds still vary in which rule they apply. Finally, similar to Sec. 3.3, we confirm that the bi-modality is still observed under continuous metric in Figure 9 and is statistically significant under Hartigan’s Dip Test (Hartigan and Hartigan, 1985).
@@ 5 Discussion
Our work explores the evolution of random variation in model performance across scales, bringing a nuanced perspective on emergent capabilities. While the mode of “emergent” performance distributions may sharply improve at a certain model scale, we attribute these sudden jumps to gradual improvements in the random distribution. In fact, bimodality often emerges before the mode—or most individual runs—exhibits a breakthrough, even when the transition from a unimodal distribution to a bimodal one is sudden.
Random variation:
@@
Model performance can be sensitive to stochastic aspects of the training process like random initialization and training data order. While variation may have a benign effect on in-distribution performance (Jordan, 2024), previous works indicate that out-of-distribution performance can fluctuate substantially across training runs, even at larger scales (Madaan et al., 2024). Prior studies have documented performance differences across various stress test sets (D’Amour et al., 2022; Naik et al., 2018), including length generalization (Zhou et al., 2024b, a). More generally, out-of-distribution behavior like compositional rules (McCoy et al., 2019) or associative biases (Sellam et al., 2022) often exhibit extreme variation compared to in-distribution loss. Such differences persist throughout training, not just at the final checkpoint (Zhou et al., 2020). Dodge et al. (2020) compared the impacts of weight initialization and data ordering, concluding that both contribute equally to variation in performance.
Existing work has also found model runs can cluster in shortcut learning (Juneja et al., 2023; Li et al., 2025) and in training dynamics (Qin et al., 2024; Hu et al., 2023), hinting at multimodal variation. Our work connects these random clustering effects to the phenomenon of emergence at scale. Relevant to our setting, Zhou et al. (2024a) and Zhou et al. (2024b) provided evidence of variability in length generalization across random seeds, which we further analyze across model scales. We also expand on the findings of Qin et al. (2024) of bimodal variation in LM syntax acquisition.
Emergent abilities of LMs:
@@
In LMs, scaling laws predict reliable performance improvements as models increase in size or train on larger datasets (Hestness et al., 2017; Rosenfeld et al., 2019; Brown et al., 2020; Kaplan et al., 2020). Emergent abilities are abilities that arise unexpectedly, out of line with these predictions (Srivastava et al., 2023; Ganguli et al., 2022; Wei et al., 2022). These abilities are characterized by unpredictable and abrupt performance improvements on specific benchmarks at certain scales. Although recent studies suggest that some breakthroughs may stem from the choice of evaluation metrics rather than fundamental changes in model behavior (Srivastava et al., 2023; Schaeffer et al., 2024), other breakthrough capabilities remain emergent—8% of high-breakthroughness datasets in BIG-Bench, according to Schaeffer et al. (2024). Our findings suggest that careful statistical analysis is required to evaluate claims both for and against “emergence” scaling on specific tasks.
Snell et al. (2024) found that some scales exhibit earlier emergence if finetuned explicitly on an emergent task, suggesting that smaller models may have the capacity for that task but are limited by its scarcity in the training corpus. Similarly, we show that emergent capabilities can arise from multimodal random variation using synthetic length generalization tasks as a case study.
Depth versus Width Scaling:
@@
Downstream performance varies with architecture shape, not just model size (Tay et al., 2022). For compositional tasks, deeper models often generalize better, but for fixed compute budgets, shallower and wider models may be advantageous (Petty et al., 2024). Several explanations of this benefit have been proposed. Edelman et al. (2024) argued that wider networks offer more parallel queries over randomized subnetworks and therefore learn sparse features efficiently. Levine et al. (2020) established width-dependent depth thresholds beyond which depth yields diminishing returns (Levine et al., 2020). We investigate how independently scaling width and depth influences random variation in compositional tasks, documenting a surprising regime (Appx. C.2) where increasing width damages performance while increasing depth improves it. Our findings should inspire further study of how emergent tasks respond to architectural hyperparameter tradeoffs.
@@ Appendix A Experimental details on synthetic tasks
Below we provide more details on the count and addition task settings. Hyperparameters for our decoder-only Transformer models are largely adapted from (Zhou et al., 2024a). We train all of our models to convergence on the train distribution. Our experiments were run on an internal cluster and all model scales can be run on a single 40GB A100 GPU with gradient accumulation; for count, runs can finish within 2 hours and for addition, runs can finish within 6 hours. At each scale, we train models from 250 seeds for the count task and 200 seeds for reverse order addition.
Count task: For all of our training runs, we fix the vocabulary size to 150. For thresholded evaluation, we compute the exact match (EM) accuracy across all consecutive subsequences of the test length.
  • Model scales: As mentioned in Sec. 2.1, we scale up our models by fixing width and scaling depth and fixing depth and scaling width. The precise parameters for each variation are as follows. For our fixed depth experiments, we fix the network depth to 4 layers and vary width by taking hidden dimensions $\{64,128,256,384,512,640,768,1024\}$. The head dimension is fixed to 64. For our fixed width experiments, we fix the hidden dimension to be 512 and vary the depth from $\{1,2,4,6,8\}$ layers.
@@
Reverse Order Addition with Index Hints: For thresholded evaluation, we compute the exact match (EM) accuracy across 500 batches of 128 examples each.